OpenStudy (anonymous):

I need some help with Algebra please! What is the simplified form of each expression? 1. –(10)–1 (1 point) -1/10 -1/1 10 1/10 10 2. 1/c -5 (1 point) c5 5c –c5 – 5/ c 3. What is the value of y-5/x -3 for x = 2 and y = –4? (1 point) 10/3 128 –1/128 –128

OpenStudy (anonymous):

I'll medal whoever can help me out i need to turn this in asap and i am absolutely terrible at math but cant let my grade get any lower, anyone that goes to indiana connections help me out please!

OpenStudy (anonymous):

I know number one is 10. I don't understand the others though.

zepdrix (zepdrix):

\[\large\rm -(10)^{-1}\]Is this the first one? :o

zepdrix (zepdrix):

Recall that for negative exponents we have this going on\[\large\rm x^{-a}=\frac{1}{x^a}\]And,\[\large\rm \frac{1}{x^{-b}}=x^b\]

OpenStudy (anonymous):

if it is then 10 won't be the answer it would be -1/10

OpenStudy (anonymous):

how are you doing that? i couldn't put the questions in correctly thats why there kind of hard to figure out what im asking

OpenStudy (anonymous):

but yes that is the first one

zepdrix (zepdrix):

The equation tool. It takes a little getting used to though :)

OpenStudy (anonymous):

\[\frac{ 1 }{ c } ^{-5}\] well then this would be question #2 but the exponet on the bottom not the top lol i attempted to use the equation tool

zepdrix (zepdrix):

So use the second rule I posted :)

zepdrix (zepdrix):

Apply the exponent rule to your 10, flip it in order to change the exponent to positive.\[\large\rm -(10)^{-1}=-\frac{1}{(10)^{+1}}=-\frac{1}{10}\]You were fine with the first one?

OpenStudy (anonymous):

Oh okay thank you, would number 2 be -c5?

zepdrix (zepdrix):

To get rid of the negative on the exponent, flip the c up to the numerator,\[\large\rm \frac{1}{c^{-5}}=\frac{c^{+5}}{1}=c^5\]Be careful with this though, I'm NOT exchanging the c and 1. I'm only moving the c. Example, if this had been your problem:\[\large\rm \frac{2}{c^{-5}}\]It would become\[\large\rm \frac{2c^{+5}}{1}=2c^5\]You're just flipping the c up to the numerator, and the exponent changes to positive. All that's left in the bottom is a 1 pretty much.

OpenStudy (anonymous):

So it would be \[c^{5}\] not \[-c^{5}\]

zepdrix (zepdrix):

correct :)

OpenStudy (anonymous):

Thank you so much! Do you think you can help me with number 3 as well?

zepdrix (zepdrix):

Apply your exponent rule to each of these BEFORE you plug the values in.